OpenStudy (anonymous):

Given the triangle and equation below, what is the value of x? Right triangle ABC with AB measuring 5 and BC measuring 7. tan∠A - csc∠C = 3x

OpenStudy (anonymous):



OpenStudy (phi):

First find length AC using pythagorean theorem. tan A = opp/adj = 7/5 csc C = 1/sinC where sin C = 5/hypotenuse (see above, AC) solve for x

OpenStudy (anonymous):

hypotenuse is sqrt of 74, but im so confused can u please solve?

OpenStudy (phi):

so what is sin(C) ? what is 1/sin(C) ? hint if sin(C)= a/b 1/sin(C)= b/a

OpenStudy (anonymous):

what?????????????????????????????? can youy please just give me an answer??? i dont understand,

OpenStudy (phi):

Watch this if you need more info http://www.khanacademy.org/video/basic-trigonometry?playlist=Trigonometry

OpenStudy (anonymous):

omgg....b being what??? and a being what?????